OpenStudy (waleed_imtiaz):

integral of cos x / (2 - cos x) dx

OpenStudy (waleed_imtiaz):

please give me the complete solution .

OpenStudy (hba):

\[\int\limits_{}^{} \cos x / (2 - \cos x) dx\]

OpenStudy (anonymous):

\[ {\cos x \over 2 - \cos x} = {}\left( {2\over 2 -\cos x} - 1 \right)\] now use Weierstrass substitution

OpenStudy (shubhamsrg):

not much need actually, after you split it as @hash.nuke did, you will find main problem is to integrate 2/(2-cosx) or 1/(1- cos x) => 1/(2 sin^2 x/2) = 1/2 * (cosec ^2 (x/2)) ,,hope rest is easier for you..

OpenStudy (sirm3d):

\[\frac{2}{2-\cos x} \neq \frac{1}{1-\cos x}\] @shubhamsrg

OpenStudy (raden):

use this : let tan(x/2) = t --> 1/2 * sec^2 (x/2) dx = dt --> dx=2*1/(t^2+1)*dt cos^2(x/2) = 1/sqrt(1+t^2) cosx = (1-t^2)/(t^2+1) so, it can be int {[(1-t^2)/(t^2+1)]/[(2 - (1-t^2)/(t^2+1]} * 2/(t^2+1)*dt = int [(2-2t^2)/((3t^2+1)(t^2+1))] dt use int by parts fraction (2-2t^2)/((3t^2+1)(t^2+1)) = A/(3t^2+1) + B/(t^2+1) gives A=4, B=-2 continue it ...... :p

OpenStudy (shubhamsrg):

ohh am so sorry! damn!!